# icme2023template.bib

### v1

@@ 1 Introduction
  • We introduce a Confidence Pruning Strategy to find the most suitable upper anchor counts of the instances, which can be used to prune and filter out the optimal anchors while taking into account the efficiency and performance, increasing the mAP by nearly 2%.
@@ 3.1 Overall Architecture
We first introduce our overall framework for action localization, where the proposed dual-stream spatiotemporal feature modeling network SVMViT is the key module as the backbone part. The framework is designed to detect all individuals in an input video clip and estimate their action labels. It is constructed using a pre-existing person detector (e.g. YOLOv9 [wang2024yolov9] ) and a video backbone network (e.g. I3D [carreira2017quo] ). The features of individuals and their surroundings are subsequently analyzed by the established ACAR module [pan2021actor] , which incorporates a long-term Actor-Context Feature Bank for the final action prediction.
@@ 4.3 Comparison with the state-of-the-art methods
Table 1: Comparison of the performance and efficiency improvements of EVAD with higher resolutions. “ft” denotes fine-tuning on this dataset. “/ ” represents unclear about its pre-training dataset.
@@ 4.4 Ablation Studies
ViT [dosovitskiy2020image] Hiera-Large K400 37.57

### v2

@@ 1 Introduction
  • We introduce a Confidence Pruning Strategy to find the most suitable anchor number of the instances, which can be used to prune anchors and filter out the optimal anchors while taking into account the efficiency and performance, increasing the mAP by nearly 2%.
@@ 3.1 Overall Architecture
We first introduce our overall framework for action localization, where the proposed dual-stream spatiotemporal feature modeling network SFMViT is the key module as the backbone part. The framework is designed to detect all individuals in an input video clip and estimate their action labels. It is constructed using a pre-existing person detector (e.g. YOLOv9 [wang2024yolov9] ) and our proposed video backbone network SFMViT. The features of individuals and their surroundings are subsequently analyzed by the established ACAR module [pan2021actor] , which incorporates a long-term Actor-Context Feature Bank for the final action prediction.
@@ 4.3 Comparison with the state-of-the-art methods
Table 1: Comparison of the performance of SFMViT with other resolutions. “ft” denotes fine-tuning on this dataset. “/ ” represents unclear about its pre-training dataset.
@@ 4.4 Ablation Studies
ViT [dosovitskiy2020image] ViT-Giant K400 37.57